Question 4 2 pts 5 years ago you purchased GDL stock for $27.6. Today, you sold the stock for $28.7. What is your annualized rate of return on this investment? Your answer must be in decimals (0.1000 is correct, NOT 10.00%). Round your answer to 4 decimals. D Question 5 2 pts You will deposit $400 per year for the next 5 years. You expect the interest rate 1 year from now to be 6%, 2 years from now to be 5%, 3 years from now to be 11%, and 4 years from now to be 6%. If your forecast of interest rates is correct, how much money will you have 5 years from now? Round your answer to 2 decimal places, for example 100.12.
Answer #1

Question 4

Capital gain = sale price-purchase price=$28.7-$27.6=$1.1

The gain has been achieved over 5 years; so annual return = CAGR = {(final value/initial value)^(1/no of years)}-1={(28.7/27.6)^(1/5)}-1=0.0078

Question 5

Assumption - compound interest is used

Amount = principal*(1+ interest rate)

Year Principal at start of year $ Interest rate Amount at end of year $ Amount calculation
1 400 no value provided in question; assumed to be 6% 424 =400*(1+6%)
2 =424+400 6% 873.44 =824*(1+6%)
3 =873.44.2+400 5% 1337.11 =1273.44*(1+5%)
4 =1337.11+400 11% =1928.19 =1737.11*(1+11%)
5 =1928.19+400 6% =2467.88 =2328.19*(1+6%)

So answer is $2467.88